A common technique used to measure the
force constant k of a spring is the following:
Hang the spring vertically, then allow a
mass m to stretch it a distance d from the
equilibrium position under the action of the
“load” m g.Find the spring constant k if the spring
is stretched a distance 82 m by a suspended
weight of 64 N. The acceleration of gravity is
9.8 m/s2. Answer in units of N/m.
1
Expert's answer
2020-02-17T09:15:08-0500

For the equilibrium:


"kx=mg=W"

"k=\frac{W}{x}=\frac{64}{82}=0.78\frac{N}{m}"